# Layer Swapping for Zero-Shot Cross-Lingual Transfer in Large Language Models

### v1

@@ 1 Introduction
Instruction fine-tuning Large Language Models (LLMs) is necessary to customize pre-trained models for real-world applications. This fine-tuning is especially critical in multilingual settings because most popular open-source LLMs have been pretrained on highly English-centric data (Llama et al., 2024; Yang et al., 2024a; Jiang et al., 2023). While recent LLMs like Llama 3 and Qwen2 have seen over a trillion non-English tokens due to the sheer scale of their pretraining datasets, these tokens are still heavily concentrated in just a few languages, resulting in limited capabilities for most other non-English languages (Llama et al., 2024; Yang et al., 2024a). Furthermore, the scarcity of high-quality labeled data available for post-training in non-English languages and the tremendous cost of procuring it further exacerbates the inequality. Even machine-translating English post-training data to target languages—a typical solution—has significant computational overhead and often leads to datasets of unreliable quality (Khanuja et al., 2024). Numerous efforts have set out to annotate or assemble high-quality datasets for lower-resource languages (Khan et al., 2024; Singh et al., 2024b; Tonja et al., 2024), but a massive gap still exists in many tasks and domains, such as math. As a result, developers are forced to largely rely on cross-lingual transfer—the generalization of learned capacities from high-resource languages to lower ones—but the rate of such transfer is low for most languages (Philippy et al., 2023).
@@
In this paper, we present a novel solution that merges two LLMs together in order to transfer math reasoning capabilities to lower-resource languages during supervised fine-tuning (SFT). In the absence of in-language math data, we fine-tune two variants of the same pretrained model: one with English math samples and the other with generic instruction data in the target language. Provided these variants, which we refer to as “experts”, we combine their learned language and task capabilities by re-composing an LLM with a mix of parameters from each. Notably, the top and bottom few transformer layers are selected from the language expert and the middle transformer layers are selected from the math expert. In addition, we build 1- or 2-layer “transition” zones in between, which consists of weighted averages of the parameter values of both. This simple, yet strategic, recomposition leads to high performance in math in the target language and avoids negative interference. The intuition behind this layer swapping method is informed by our analysis of the SFT updates that led to the experts. We find that the learned math capabilities are concentrated in the middle transformer layers, especially in the second half of the model. Meanwhile, we find the enhanced language skills to have come from parameters closest to the input and output, in line with previous literature stating this is where the most language-specific representations are concentrated (Chang et al., 2022; Choenni et al., 2024). Therefore, our methodology takes the most important layers from each expert and transfers math capabilities to the target language.
The resulting layer-swapped models deliver strong performance across numerous target languages (Swahili, Telugu, Bengali, Japanese) on MGSM (Shi et al., 2023), the manually-translated version of the Grade School Math benchmark (Cobbe et al., 2021). In all these major languages, where math SFT data is not readily available, layer swapping outperforms baselines, the individual experts and model souping (Wortsman et al., 2022) by 10% on average. For Swahili, layer swapping exceeds MGSM performance of models on fine-tuned a mixed Swahili and math SFT dataset when evaluated. Therefore, this methodology provides a simple and effective way to improve the capabilities of LLMs without the need for any task-specific data in the target language. In addition, it is inexpensive and fully post hoc, meaning it has the potential to be practical in many settings. Fundamentally, the success of this method also provides empirical evidence for cross-lingual patterns in the latent structures of LLMs that can be further interpreted and exploited.
@@ 2.1 Model Merging
While the use of weight averaging to stabilize training through noise reduction well pre-dates instability challenges in deep neural networks (Breiman, 1996), the use of model merging to combine trained model checkpoints is an emerging research space in deep learning. Similar to ensembling model outputs (Dietterich, 2000), aggregating model weights also improves model robustness and generalization Izmailov et al. (2018), even when combining models trained on the same data (Ramé et al., 2022). Numerous studies seek to improve the pre-merging conditions, either via linearizing fine-tuning (Ortiz-Jimenez et al., 2023) or aligning weights (Ainsworth et al., 2023). Wortsman et al. (2022) develops an empirical method to average, or soup, a subset of available model checkpoints together, in order to increase the model search space during hyperparameter tuning. Simple weight averaging, however, is at risk of negative transfer, or interference, between trained checkpoints. To address this, methods have been proposed to selectively combine models at the individual weight level, either using the magnitude and direction of the fine-tuning deltas (Ilharco et al., 2023; Yadav et al., 2023; Davari & Belilovsky, 2024; Yu et al., 2024) or leveraging information theory (Matena & Raffel, 2022). In parallel, sparse fine-tuning methods have been developed to create fine-tuned models with a small proportion of weights changed (Guo et al., 2021; Sung et al., 2021; Xu et al., 2021), which then allows adding together fine-tuning updates with less overlap. Overall, model merging, and notably inexpensive model souping, is very common in practice because it improves training stability, model robustness and generalization, and performance by increasing the search space or combining expertise in multi-task settings (Yang et al., 2024b).
@@ 2.2 LLM Multilinguality
The inability for a language model to learn more languages without undermining other capabilities—the curse of multilinguality (Conneau et al., 2020; Pfeiffer et al., 2022)—was a heavily studied problem in encoder models. Numerous investigations attempted to understand the quantity and location of parameter sharing versus language-specific parameters (Wang et al., 2020; Choenni et al., 2023) in massively multilingual encoder models (e.g. mBERT (Devlin et al., 2019)). While higher rates of language-specific parameters were found in the top and bottom layers, largely because of the proximity to token representations (Chang et al., 2022; Choenni et al., 2024), language specialization occurs across the model. In addition, increasing the model vocabulary has proven effective in increasing a model’s language capacity (Zheng et al., 2021; Liang et al., 2023; Llama et al., 2024). And while recently model scaling has mitigated the limitation from parameter quantity, the massive amount of labeled data needed for LLM training presents a new challenge. In encoder models, cross-lingual transfer was enhanced by aligning cross-lingual representations (Patra et al., 2023; Ouyang et al., 2021; Gaschi et al., 2023), but this is difficult in decoder-only models. To boost transfer in post-training, several data synthesis or augmentation solutions have been proposed, both for SFT (Zhang et al., 2023b; Qin et al., 2023; Chai et al., 2024) or reinforcement learning from human feedback (RLHF) (Dang et al., 2024; She et al., 2024; Lai et al., 2024).
@@ 2.3 Model Merging for Cross-Lingual Transfer
To address the limited representational capacity of multilingual models, many solutions have been proposed to strategically share or splits parts of the model. This could be major blocks, as in mixture-of-experts (Fedus et al., 2022; NLLB et al., 2022), or a few parameters in each layer, as in cross-lingual adapters (Pfeiffer et al., 2020; 2022). Ansell et al. (2022) combines modular and sparse fine-tuning approaches in a two-stage SFT, where first task and language experts are fine-tuned and all weights who did not change more than a threshold are masked. Next, the experts are fine-tuned again from the pretrained model with this mask, creating sparse task vectors that can be composed together with lower rates of parameter overlap. This composable sparse fine-tuning was also adapted for larger decoder-only LLMs (Ansell et al., 2024). In this work, we develop a simpler and more flexible merging method for cross-lingual transfer that does not require fine-tuning more than once.
@@ 3.1 Setup
We start by training numerous math and language “experts” by fine-tuning Llama 3.1 8B (Llama et al., 2024). We perform SFT runs with 30-40k labeled samples with varying hyperparameters [^1] Training specifics such as hyperparameters are provided in Appendix A.2 and for each type of expert, select the three best checkpoints. The math experts were fine-tuned on English math world problems from the Orca-Math synthetic dataset (Mitra et al., 2024). In order to select the three experts, we use results on the English splits of MGSM, as well as the average across languages. For the language experts, we select Swahili, Bengali, Telugu, and Japanese as the target languages. These are languages present in MGSM and other benchmarks discussed below, but where Llama 3.1’s performance lags behind the top languages like Spanish. In addition, boosting cross-lingual transfer to these languages is important because of the lack of in-language math instruction data. For each of the four, we mix together samples from available instruction datasets in that language, to create experts with enhanced general language and instruction-following capabilities. The resulting datasets contain many different types of tasks—such as translation, NER, and question-answering—but no math [^2] Dataset details are provided in Appendix A.1. After numerous SFT runs on these generic, multi-task datasets, the three checkpoints for each language are primarily selected based off their performance on the target language splits on Belebele (Bandarkar et al., 2024) and FLoRes (NLLB et al., 2022). These simple tasks, reading comprehension and translation, are strong indicators of basic language understanding and generation capabilities. We, however, also ensure slight improvement on MBPP (Austin et al., 2021), MMLU (Hendrycks et al., 2021), and MGSM as secondary measures of language improvement.
@@ 3.2 Parameter Analysis of SFT
Figure 2: This visualization displays the location with more significant magnitude of change during SFT for two representative experts. In detail, this shows the percentage of rows for each 2-dimensional parameter in the 32 transformer layers of Llama 3.1 8B where the mean absolute value is above a threshold ($1.9\times 10^{-5}$ and $1.0\times 10^{-5}$, respectively). The darker the green represents parameters changing more significantly, relative to the others. Larger versions of these images, as well as for more experts, can be found in Appendix A.4
@@ 3.3 Sparsifying Updates
We then attempt to create sparse model updates for increased composability similar to Ansell et al. (2022), but without retraining. We attempted leveraging the magnitude of the deltas to selectively update the model without undermining performance gains. Concretely, we use thresholds to determine whether to apply the update or revert the value to the original value in the pretrained Llama 3.1. This is done at row-level granularity, in order to not partially modify linear transformations. However, in our analysis across math experts, we find that more than 70-80% of model parameters are required to be updated for the increase in performance to remain equivalent. Such a small degree of sparsification would not significantly reduce interference between math and language experts.
We next attempted location-based sparsification. This means leveraging our intuition of what patterns occur during fine-tuning to select specific parameter tensors to merge or not. A breakthrough was made when we revert the first five and last two transformer layers of our math experts to their original values. After this mitigation, we find that the math experts perform equally well. This is significant because our intuition—based on our analysis and previous multilingual interpretability research on both encoder-decoder models (Chang et al., 2022; Choenni et al., 2024) and decoder-only models (Tang et al., 2024; Zhang et al., 2024b)—is that the first few and last few transformer layers contain the most important language-specific parameters. Meanwhile, our discovery implies the first few and last few transformer layers were not very important for the math experts. This contrast prompted the swapping of these layers from the language expert into the math expert.
@@ 4.1 Layer Swapping
The layer swapping methodology takes two fine-tuned experts, one on the target language and the other on the target task in English, and re-composes a joint model with the top and bottom transformer layers from the language expert and the middle from the math expert.
As displayed in Fig. 1, we additionally design a transition zone in between the off-the-shelf layers from each expert. These buffer layers are weighted averages of the respective layers from each expert that ensure that the outputs of one layer does not directly input into a layer fine-tuned separately. While intuitively these transition zones seem necessary, we do not find empirical evidence that they provide statistically significant benefit over replacing them with the math expert’s layers. However, as discussed in Section 3.1, our experts were not fine-tuned for very long and therefore the latent representation spaces after each layer would not have diverged very much from each other. This explains, in theory, why transition zones were not helpful in our setting. We conjecture that if the experts had been trained further (or simply with higher learning rates), such a buffer zone would be necessary. We therefore still present it as a central component of our layer swapping methodology.
@@
0:      task expert $\theta_{task}$, language expert $\theta_{lang}$, lower layers to swap $b$, upper layers to swap $u$, lower transition layers $t_{b}$, upper transition layers $t_{u}$, weight of each expert $w_{task}$, $w_{lang}$, number of model layers $L$
@@
4:        $\theta_{merged}\{n\}\leftarrow\theta_{lang}\{n\}$
@@
6:        $\theta_{merged}\{n\}\leftarrow\theta_{task}\{n\}$
@@
8:        $\theta_{merged}\{n\}\leftarrow(w_{task}*\theta_{task}\{n\}+w_{lang}*\theta_{lang}\{n\})/(w_{task}+w_{lang})$
@@ 4.2 Configuration
    The transition zones are averages of the layers from both (i.e. soups) that can be unweighted or magnitude-adjusted weighted averages.
@@
    The non-transformer parameters (input token embeddings, output layers, etc) work best if they are also averages of the two experts, as opposed to simply from the language expert.
@@
While the results were largely equivalent within this range, some patterns did exist. We find that for languages where performance was lower (e.g. Telugu), the higher performing configurations typically had more layers from language expert. This is perhaps because the boost from improving the language capabilities was more important relative to languages Llama 3.1 was already better at. For example for languages such as Japanese, the best configuration has close to the minimum layers from language expert, similar to the minimum swap illustrated in Fig. 3. Amongst the math experts, the math expert with the highest scores on all four target languages (named “math expert #2”) performed understandably better with fewer layers from the language expert swapped in. Generally, we conjecture that lower-resource languages tend to require more layers from the language expert.

### v2

@@ 1 Introduction
Instruction fine-tuning Large Language Models (LLMs) is necessary to customize pre-trained models for real-world applications. This fine-tuning is especially critical in multilingual settings because most popular open-source LLMs have been pretrained on highly English-centric data (Jiang et al., 2023; Llama et al., 2024; Yang et al., 2024a). Although recent LLMs such as Llama 3 and Qwen2 have seen more than a trillion non-English tokens due to the sheer scale of their pretraining datasets, these tokens are still heavily concentrated in just a few languages, resulting in limited capabilities for most other non-English languages (Llama et al., 2024; Yang et al., 2024a). Furthermore, the scarcity of high-quality labeled data available for post-training in non-English languages and the tremendous cost of procuring it further exacerbates the inequality. Even machine-translating English post-training data into target languages—a typical solution—has significant computational overhead and often leads to datasets of unreliable quality (Khanuja et al., 2024). Numerous efforts have set out to annotate or assemble high-quality datasets for lower-resource languages (Khan et al., 2024; Singh et al., 2024b; Tonja et al., 2024), but a massive gap still exists in many tasks and domains, such as math. As a result, developers are forced to largely rely on cross-lingual transfer—the generalization of learned capacities from high-resource languages to lower ones—but the rate of such transfer is low for most languages (Philippy et al., 2023).
@@
In this paper, we present a novel solution that merges two LLMs together in order to transfer math reasoning capabilities to lower-resource languages during supervised fine-tuning (SFT). In the absence of in-language math data, we fine-tune two variants of the same pretrained model: one with English math samples and the other with generic instruction data in the target language. Provided these variants, which we refer to as “experts”, we combine their learned language and task capabilities by re-composing an LLM with a mix of parameters from each while avoiding negative interference. Notably, the top and bottom few transformer layers are selected from the language expert, and the middle transformer layers are selected from the math expert. The intuition behind this simple, yet strategic, method is informed by our analysis of the SFT updates that led to the experts. We find that the learned math capabilities are concentrated in the middle transformer layers, especially in the second half of the model. Meanwhile, we find the enhanced language skills to have come from parameters closest to the input and output, in line with previous literature stating that this is where the most language-specific representations are concentrated (Wendler et al., 2024; Alabi et al., 2024; Tang et al., 2024). Therefore, our methodology takes the most important layers from each expert and transfers math capabilities to the target language.
The resulting layer-swapped models deliver strong performance across numerous target languages (Swahili, Telugu, Bengali, Japanese) on MGSM (Shi et al., 2023), the manually-translated version of the Grade School Math benchmark (Cobbe et al., 2021). In all these major languages, where math SFT data is not readily available, layer swapping outperforms baselines, the individual experts, and model souping (Wortsman et al., 2022)) by 10% on average. For Swahili, layer swapping exceeds MGSM performance of models fine-tuned on a mixed Swahili and math SFT dataset when evaluated. Therefore, this methodology provides a simple and effective way to improve the capabilities of LLMs without the need for any task-specific data in the target language. In addition, it is inexpensive and fully post hoc, meaning that it has the potential to be practical in many settings. Fundamentally, the success of this method also provides empirical evidence for cross-lingual patterns in the latent structures of LLMs that can be further interpreted and exploited.
@@ 2.1 Model Merging
While the use of weight averaging to reduce noise in training well predates instability challenges in deep neural networks (Breiman, 1996), the use of model merging to combine trained model checkpoints is an emerging research space in deep learning. Similar to ensembling model outputs (Dietterich, 2000), aggregating model weights also improves model robustness and generalization (Izmailov et al., 2018), even if trained on the same data (Ramé et al., 2022). Numerous studies seek to improve the pre-merging conditions, either via linearizing fine-tuning (Ortiz-Jimenez et al., 2023) or aligning weights (Ainsworth et al., 2023). Wortsman et al. (2022) develops an empirical method to average, or soup, model variants together to increase the search space during hyperparameter tuning. However, simple weight averaging, is vulnerable to negative transfer, or interference, between variants. To address this, methods have been proposed to selectively combine models at the individual weight level, using either the magnitude and direction of the fine-tuning deltas (Ilharco et al., 2023; Yadav et al., 2023; Davari & Belilovsky, 2024; Yu et al., 2024) or leveraging information theory (Matena & Raffel, 2022). In parallel, sparse fine-tuning methods have been developed to create fine-tuned models with a small proportion of weights changed (Guo et al., 2021; Sung et al., 2021; Xu et al., 2021), which then allows adding together fine-tuning updates with less overlap. Overall, model merging, especially inexpensive model souping, is very common in practice because it improves training stability, model robustness and generalization, and performance by increasing the search space or combining expertise in multi-task settings (Yang et al., 2024b).
@@ 2.2 LLM Multilinguality
The inability for a language model to learn more languages without undermining other capabilities—the curse of multilinguality (Conneau et al., 2020; Pfeiffer et al., 2022)—was a heavily studied problem in encoder models. Beyond increasing vocabulary capacity (Zheng et al., 2021; Liang et al., 2023), numerous works attempt to understand the quantity and location of language-specific parameters (Wang et al., 2020; Muller et al., 2021; Choenni et al., 2023) in multilingual encoders. While higher rates of language-specific parameters were found in the top and bottom layers (Chang et al., 2022; Choenni et al., 2024), language specialization occurs throughout the model. And while recently model scaling has mitigated the limitation from parameter quantity, the massive amount of labeled data needed to train LLMs presents a new challenge. In encoder models, cross-lingual transfer was enhanced by aligning cross-lingual representations (Ouyang et al., 2021; Patra et al., 2023; Gaschi et al., 2023), but this is difficult in decoder-only models (Jain et al., 2023). To boost transfer in post-training, several data augmentation solutions have been proposed, both for SFT (Qin et al., 2023; Chai et al., 2024) or reinforcement learning from human feedback (RLHF) (Dang et al., 2024; She et al., 2024; Lai et al., 2024). Recent work finds that English-centric LLMs process multilingual text by mapping them to and from English representations in the first and last transformer layers to take advantage of English-based capabilities (Kojima et al., 2024; Wendler et al., 2024; Tang et al., 2024; Alabi et al., 2024). This leads to benefits from prompting such LLMs to “think in English”, thereby promoting this internal mapping (Shi et al., 2023; Zhang et al., 2024b).
@@ 2.3 Model Merging for Cross-Lingual Transfer
To address the limited representational capacity of multilingual models, many solutions have been proposed to strategically share or split parts of the model. This could be major blocks, as in mixture-of-experts (Fedus et al., 2022; NLLB et al., 2022), or a few parameters in each layer, as in cross-lingual adapters (Pfeiffer et al., 2020; 2022). Ansell et al. (2022) combines modular and sparse fine-tuning approaches in a two-stage SFT, where first task and language experts are fine-tuned and all weights that did not change more than a threshold are masked. Next, the experts are fine-tuned again from the pretrained model with this mask, creating sparse task vectors that can be composed together with lower rates of parameter overlap. This composable sparse fine-tuning was also adapted for larger decoder-only LLMs (Ansell et al., 2024). In this work, we develop a simpler and more flexible merging method for cross-lingual transfer that does not require fine-tuning more than once.
@@ 3.1 Setup
We start by training numerous math and language “experts” by fine-tuning Llama 3.1 8B (Llama et al., 2024). We perform SFT runs using next token prediction with 30-40k labeled samples with varying hyperparameters [^1] Training specifics such as hyperparameters are provided in Appendix A.2 and for each type of expert, select the three best checkpoints. The math experts were fine-tuned on English math word problems from the Orca-Math synthetic dataset (Mitra et al., 2024). In order to select the three experts, we use results on the English splits of MGSM, as well as the average across languages. For the language experts, we select Swahili, Bengali, Telugu, and Japanese as the target languages. These are languages present in MGSM and other benchmarks discussed below, but where Llama 3.1’s performance lags behind the top languages like Spanish. In addition, the lack of in-language math instruction data in these languages motivates the need for effective methods for zero-shot cross-lingual transfer. For each, we mix together samples from available instruction datasets in that language to create experts with enhanced general language and instruction-following capabilities. The resulting datasets contain many different types of tasks—such as translation, NER, and question-answering—but no math [^2] Dataset details are provided in Appendix A.1. After numerous SFT runs on these “generic” multi-task datasets, the three checkpoints for each language are primarily selected based off their performance on the target language splits on Belebele (Bandarkar et al., 2024) and FLoRes (NLLB et al., 2022). These simple tasks, reading comprehension and translation, are strong indicators of basic language understanding and generation capabilities. We also ensure slight improvement on MBPP (Austin et al., 2021), MMLU (Hendrycks et al., 2021), and MGSM as secondary measures of language improvement  [^3] See results of chosen experts in Appendix A.3.
@@ 3.2 Parameter Analysis of SFT
Figure 2: This visualization displays the location with more significant magnitude of change during SFT for two representative experts across transformer layers and parameter types. In detail, this shows the percentage of rows for each 2-dimensional parameter in the 32 transformer layers of Llama 3.1 8B where the mean absolute value is above a threshold ($1.9\times 10^{-5}$ and $1.0\times 10^{-5}$, respectively). Darker green shading represents parameters changing more significantly, relative to the others. Larger versions of these images, as well as for more experts, can be found in Appendix A.5
@@ 3.3 Sparsifying Updates
We then attempt to create sparse model updates for increased composability similar to Ansell et al. (2022), but without retraining. We attempt to utilize the magnitude of the deltas to selectively update the model without undermining performance gains. Concretely, we use thresholds to determine whether to apply the update or revert the value to the original value in the pretrained Llama 3.1. This is done at row-level granularity in order to not partially modify linear transformations. However, in our analysis across the math experts, we find that more than 70-80% of model parameters are required to be updated for the increase in performance to remain equivalent. Such a small rate of sparsification would not significantly reduce interference between the math and language experts.
We next attempt location-based sparsification. This means leveraging our intuition of what patterns occur during fine-tuning to select specific parameter tensors to merge or not. A breakthrough was made when we revert the first five and last two transformer layers of our math experts to their original values. Despite undoing the SFT updates, the math experts maintain their increased performance. This is significant when considered alongside our intuition that the first few and last few transformer layers contain the most important language-specific parameters. This is based on our above analysis and is supported by previous multilingual interpretability research on both encoder-decoder models (Chang et al., 2022; Choenni et al., 2024) and decoder-only models (Tang et al., 2024; Zhang et al., 2024c). Our discovery that SFT updates to these same layers are not critical for mathematical reasoning led us to experiment with swapping them in from the language expert.
@@ 4.1 Layer Swapping
The layer swapping methodology takes two experts, one fine-tuned on the target language and the other on the target task in English, and re-composes a single model with the top and bottom transformer layers from the language expert and the middle from the math expert.
As displayed in Figure 1, we additionally design a transition zone in between the off-the-shelf layers from each expert. These buffer layers are weighted averages of the respective layers from each expert that ensure that the output of one layer does not directly input into a layer fine-tuned separately. While intuitively these transition zones seem necessary, we do not find empirical evidence that they provide statistically significant benefit over replacing them with the math expert’s layers. However, as discussed in Section 3.1, our experts were not fine-tuned for very long and therefore the latent representation spaces after each layer would not have diverged very much from each other. This explains, in theory, why transition zones were not helpful in our setting. We conjecture that if the experts had been trained further (or simply with higher learning rates), such a buffer zone would be necessary. We therefore still present it as a central component of our layer swapping methodology.
@@
0:      task expert $\theta_{task}$, language expert $\theta_{lang}$, lower layers to swap $b$, upper layers to swap $u$, lower transition layers $t_{b}$, upper transition layers $t_{u}$, weight of each expert $\alpha_{task}$, $\alpha_{lang}$, number of model layers $L$
@@
4:        $\theta_{merged}\{n\}\leftarrow\theta_{lang}\{n\}$ top & bottom layers from language expert
@@
6:        $\theta_{merged}\{n\}\leftarrow\theta_{task}\{n\}$ middle layers from task expert
@@
8:        $\theta_{merged}\{n\}\leftarrow(\alpha_{task}*\theta_{task}\{n\}+\alpha_{lang}*\theta_{lang}\{n\})/(\alpha_{task}+\alpha_{lang})$ transition zone
@@ 4.2 Configuration
    The transition zones are averages of the layers from both (i.e. soups) that can be unweighted or magnitude-adjusted weighted averages ($\alpha_{task},\alpha_{lang}$).
@@
    The non-transformer parameters (input token embeddings, output layers, etc.) work best if they are also averages of the two experts, as opposed to simply from the language expert.
@@
Although the results were largely equivalent within this range, some patterns did exist. We find that for languages where performance was lower (e.g. Telugu), the higher performing configurations typically had more layers from the language expert. This is perhaps because the boost from improving the language capabilities was more important relative to languages Llama 3.1 was already better at. For example for languages such as Japanese, the best configuration has close to the minimum layers from language expert, similar to the minimum swap illustrated in Figure 3. The math expert where the most cross-lingual transfer occurred naturally during SFT (that is, with the highest scores on all four target languages), named “math expert #2”, performed understandably better with fewer layers from the language expert swapped in. Generally, we conjecture that lower-resource languages tend to require more layers from the language expert.